OpenStudy (anonymous):

The graph reaches it's maximum when x=2. What is the y value of this maximum?

OpenStudy (anonymous):

What is the equation of the graph?

OpenStudy (anonymous):

y=-x^2+4x

OpenStudy (anonymous):

You just have to plug in.

Directrix (directrix):

y=-x^2+4x y = - ( 2 ) ^ 2 + 4 ( 2 ) y = ?